Landforms Produced by Running Water and Groundwate Promontory Butte Quadrangle, Arizona a. Trace in a red pencil the drainage divide betwe b. What human-made feature is roughly co-incident with the divide? c. The table below lists several streams north (A) and south (B) of the Mogollon R en the sources of the two stream systems im. For each stream, the differences in elevation of the two contour the map distance between them is shown. lines that cross the stream a few miles apart are recorded, and 1. For each stream segm ent listed in the table, calculate its gradient in feet per mile 2. Determine the average gradient for the streams north and south of the Mogollon kim 3. If the erosive potential of a stream is proportional to its gradient, which system, A the greatest erosive power? Will the div the passage of time? Explain your answer 4. ide between the two systems move generally north or south, or remain the same with
the greatest erosive power? Will the divide between the two systems move generally north or sou the passage of time? Explain your answer th, c 4. TABLE1-PROMONTORY BUTTE GRADIENTS A. Group- Stream gradients north of Mogollon Rim Gradient Name of Stream West Leonard Canyon Middle Leonard Canyon East Leonard Canyon Turkey Canyon Beaver Canyon Bear Canyon Length in Miles Diff. Elev. Ft. 2.6 2.4 1.9 1.9 3.7 1.6 250 400 250 250 300 200 B. Group - Stream gradients south of Mogollon Rim Name of Stream Big Canyon Creek Dick Williams Creek Horizon Creek Doubtful Canyon Spring Creek See Canyon Length in Miles Diff. Elev. Ft. 3.7 1.8 3.0 2.4 2.2 Gradient 1,500 1,150' 1,000 1.250 1,250 700 2.2 Page 2 of 13
Answer #1

a. If the quadrangle map was provided, the marking the drainage divide with red pencil could be done !

b. In general, DAMs are the human made features , which are roughly coincident with the divide.

c. 1) Calculation of stream gradient (in feet/mile):

Group A :

  • West leonard canyon = 250/2.6 = 96.1538
  • Middle leonard canyon = 400/2.4 = 166.6667
  • East leonard canyon = 250/1.9 = 131.5789
  • Turkey canyon = 250/1.9 = 131.5789
  • Beaver canyon = 300/3.7 = 81.0811
  • Bear canyon = 200/1.6 = 125

Group B :

  • Big canyon creek = 1500/3.7 = 405.4054
  • Dik williams creek = 1150/1.8 = 638.8889
  • Horizon creek = 1000/3 = 333.3333
  • Doubtful canyon = 1250/2.4 = 520.8333
  • Spring creek = 1250/2.2 = 568.1818
  • See canyon = 700/2.2 = 318.1818

2) Average gradient of streams north of Mogollon Rim (Group A streams) = (96.1538 + 166.6667 + 131.5789 + 131.5789 + 81.0811 + 125)/ 6 = 732.0594/6 = 122.0099 feet/mile.

Average gradient of streams south of Mogollon Rim (Group B streams) = (405.4054 + 638.8889 + 333.3333 + 520.8333 + 568.1818 + 318.1818)/6 = 2784.8245/6 = 464.1374 feet/mile.

3) As the erosive potential of the stream is proportional to its gradient, Group B would have more erosive power because its avg. gradient is more (464.1374 feet/mile) than the avg. gradient of the Group A streams (122.0099 feet/mile).